[Brahms-l] Shift report 20050128 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Jan 28 2005 - 08:23:33 EST
1. Problems in the beginning of shift. They went down to 37 bunches again, and got a store from ~4am. Nice low background.

2. run trig2 trig3       
12548   122k   83k
12549   110k   75k
12550    95k   66k

3. RunError when i started shift but easily fixed with resart of ioserver and runserver.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jan 28 08:24:15 2005

This archive was generated by hypermail 2.1.8 : Fri Jan 28 2005 - 08:24:28 EST